OpenStudy (anonymous):

how do i simplify this 3.25(-12.5-2y)/(-4.5)-y = -0.75

OpenStudy (anonymous):

combine like terms and solve for y.

OpenStudy (anonymous):

Do you know how to start??

OpenStudy (anonymous):

so 3.25(-12.5-3y)/(-4.5)=-0.75 whats next

OpenStudy (anonymous):

divide 3.25 on both sides.

OpenStudy (anonymous):

You'll get (-12.5-3y)/(-4.5)=-.75/3.25

OpenStudy (anonymous):

so-12.5y/-4.5=.02307

OpenStudy (anonymous):

Sorry, let me restate. Distribute the 3.25 to the (-12.5-3y).

OpenStudy (anonymous):

You get -(0.7222(-2y-12.5))-y=-.75

OpenStudy (anonymous):

Eventually you'll get y=-22

OpenStudy (anonymous):

I can explain the process step by step, but it would be long. What you really have to do is do the opposite of what the sign shows. Say a value is added to another value. What you do is subtract that value from both sides. If the values is a fraction, say 7/3y=9, what you do is multiply both sides by 7 so you'll have 3y=63.

OpenStudy (anonymous):

now you have 3y=63, Since you want to find out what y is, all what you do is do the opposite of y. Y is currently multiplied by 3, so what you do is divide both sides by 3.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I hope you get the point.

OpenStudy (anonymous):

well i really should restate the question because im trying to use the substitution method to solve a system of equations and the two equations r in the attachments

OpenStudy (anonymous):

im reposting the question

OpenStudy (anonymous):

Wait, so what is the original question?

OpenStudy (anonymous):

i just reposted it

OpenStudy (anonymous):

Is this a different question?

OpenStudy (anonymous):

Because the first question had no x value in it.